Prove that (n+1)2 n for all n>1.

Answer by rothauserc(4718)

if n = k+1 (k > 1), we have that
:
(k+1) 2(k+1)
:
we have to prove that
:
(k+2) 2(k+2)
:
(k+2)! = (k+2)(k+1) (k+2)2(k+1) (since (k+1) 2(k+1))
:
this implies that
:
(k+2) 2(k+2)